### v4

# Adapting Large Language Models for Document-Level Machine Translation

Abstract
Large language models (LLMs) have significantly advanced various natural language processing (NLP) tasks. Recent research indicates that moderately-sized LLMs often outperform larger ones after task-specific fine-tuning. This study focuses on adapting LLMs for document-level machine translation (DocMT) for specific language pairs. We first investigate the impact of prompt strategies on translation performance and then conduct extensive experiments using two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our results show that specialized models can sometimes surpass GPT-4 in translation performance but still face issues like off-target translation due to error propagation in decoding. We provide an in-depth analysis of these LLMs tailored for DocMT, examining translation errors, discourse phenomena, strategies for training and inference, the data efficiency of parallel documents, recent test set evaluations, and zero-shot crosslingual transfer. Our findings highlight the strengths and limitations of LLM-based DocMT models and provide a foundation for future research.
Adapting Large Language Models for Document-Level Machine Translation
[…]
1 Introduction
Large language models (LLMs) demonstrate impressive proficiency in a wide range of applications (Ouyang et al., 2022; Wei et al., 2022a; Sanh et al., 2022; Chung et al., 2022; OpenAI, 2023; Anil et al., 2023; Touvron et al., 2023a, b; Jiang et al., 2023). However, in the realm of translation tasks, only few very large models, such as gpt-3.5-turbo and gpt-4-turbo, can match or surpass the performance of state-of-the-art supervised encoder-decoder models like NLLB (Costa-jussà et al., 2022), while they still under-perform in translating low-resource languages (Robinson et al., 2023; Jiao et al., 2023; Hendy et al., 2023). Consequently, a number of recent works attempt to bridge the gap between LLMs and supervised encoder-decoder models in translation tasks (Zhu et al., 2023; Yang et al., 2023; Zhang et al., 2023; Moslem et al., 2023; Xu et al., 2023; Kudugunta et al., 2023). Recently, research suggests that smaller, specialized models can outperform larger, general-purpose models in specific tasks (Gunasekar et al., 2023; Luo et al., 2023; Azerbayev et al., 2023). Therefore, we explore adapting LLMs for document-level machine translation (DocMT) in this study.
In this study, we explore the capabilities of moderately-sized Large Language Models (LLMs) with 7 billion parameters across 18 translation tasks involving nine language pairs. We conduct experiments under both bilingual and multilingual settings, employing Parameter-Efficient Fine-Tuning (PEFT) and Fully Fine-Tuning (FFT) techniques on three LLM backbones. Our findings reveal that, after fine-tuning, these LLMs demonstrate superior translation performance compared to state-of-the-art models, as evidenced by metrics such as $s$BLEU, $d$BLEU, and COMET. However, a significant challenge identified is the issue of off-target translations, which persists even after exclusive fine-tuning on bilingual corpora. Our analysis attributes this high rate of off-target translations to error propagation during the inference. Additionally, we provide a comprehensive analysis of our LLM-based DocMT models from various perspectives. This includes examining translation error distribution, discourse phenomena, training strategies, the data efficiency of parallel documents, and additional evaluations on the recent WMT2023 test sets. We also investigate zero-shot cross-lingual transfer, aiming to enhance the understanding and efficacy of LLMs in DocMT tasks.
We present extensive empirical evidence that highlights both the superior translation capabilities and limitations of the LLM-based DocMT models in this study, making several significant discoveries. Here are the main takeaways:
  • Selective Excellence in Translation Tasks: Our findings show that our moderately-sized LLMs outperform gpt-4-turbo in certain translation tasks, but struggle in others due to the off-target translation issue due to error propagation in decoding. Despite this, our DocMT models exhibit better context awareness and fewer errors, while maintaining comparable performance.
  • Fine-Tuning Strategies: Our research indicates that the PEFT approach outperforms the FFT approach overall. However, the FFT approach shows greater data efficiency, needing only about $1\%$ of the total dataset to reach the performance level of models trained on the entire dataset. In contrast, the PEFT approach requires $10\%$ of the total dataset for comparable results.
  • Evaluation on Recent Test Sets: We evaluate our models on recent test sets between English and German from WMT2023 (Koehn et al., 2023). Our empirical results show that, when the data leakage risks are mitigated, the LLM-based DocMT models generalize better on out-of-domain text, compared to the conventional DocMT models.
  • Advantage of Base LLMs for Task-Specific Supervised Fine-Tuning: Our study shows that base LLMs, when used as the backbone for task-specific supervised fine-tuning, perform better than instruction-tuned LLMs. They demonstrate more effective zero-shot cross-lingual transfer.
2 Related Work
Document-Level Machine Translation
In recent years, numerous approaches have been proposed for document-level machine translation (DocMT). There exist other approaches to DocMT, including document embedding (Macé and Servan, 2019; Huo et al., 2020), multiple encoders (Wang et al., 2017; Bawden et al., 2018; Voita et al., 2018; Zhang et al., 2018), attention variations (Miculicich et al., 2018; Zhang et al., 2020; Maruf et al., 2019; Wong et al., 2020; Wu et al., 2023), and translation caches (Maruf and Haffari, 2018; Tu et al., 2018; Feng et al., 2022). Furthermore, Maruf et al. (2022) present a comprehensive survey of DocMT.
Large Language Models
Large language models (LLMs) have demonstrated remarkable proficiency across a wide range of Natural Language Processing (NLP) tasks (Brown et al., 2020; Chowdhery et al., 2022; Scao et al., 2022; Anil et al., 2023; Touvron et al., 2023a, b). Furthermore, recent research has shown that supervised fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) can significantly enhance their performance when following general language instructions Mishra et al. (2022); Wang et al. (2022); Shen et al. (2023); Li et al. (2023); Wu and Aji (2023); Wu et al. (2024a). More recently, there is a growing body of work exploring the translation capabilities of LLMs (Lu et al., 2023; Zhang et al., 2023; Xu et al., 2023; Robinson et al., 2023; Wu et al., 2024d). However, it is important to note that these efforts have primarily focused on sentence-level machine translation (SenMT) and have not delved into document-level machine translation (DocMT). A noteworthy study in DocMT is conducted by Wang et al. (2023b), where they investigate the document-level translation capabilities of gpt-3.5-turbo, making it the most closely related work to our work.
Ours
In contrast to the work of Wang et al. (2023b), who primarily investigate the use of gpt-3.5-turbo for DocMT through prompting techniques, our study concentrates on analyzing the effectiveness of parameter-efficient fine-tuning (PEFT) and full fine-tuning (FFT) methods on moderately-sized LLMs in the context of DocMT.
3 Experimental Setup
In this section, we detail our experimental setup, covering the training strategy (Section 3.1), datasets (Section 3.2), models (Section 3.3), and evaluation (Section 3.4).
3.1 Two-Stage Training
DocMT approaches typically start by pre-training the translation model on parallel sentences and then fine-tuning it on parallel documents (Voita et al., 2019; Ma et al., 2020; Wu et al., 2023). Recently, Xu et al. (2023) proposed a two-stage training strategy. This involves initially fine-tuning a LLM on monolingual text, followed by a second fine-tuning phase on parallel text. Following Xu et al. (2023), our approach begins with fine-tuning an LLM on monolingual documents, followed by fine-tuning on parallel documents.
Fine-tuning on Monolingual Documents
Existing LLMs are typically pre-trained on English-centric corpora. Recent research highlights that these LLMs often exhibit sub-optimal performance on multilingual benchmarks (Li et al., 2023; Chen et al., 2023; Scao et al., 2022). To address this limitation, our initial step involves fine-tuning all the parameters of LLMs using monolingual data from the target languages.
Fine-tuning on Parallel Documents
We fine-tune the model on document-level parallel corpora in this stage. Following Wang et al. (2023a), we condition each sentence pair on its context, consisting of the three preceding consecutive sentence pairs. As demonstrated by Wang et al. (2023b), the prompting strategy plays a significant role in translating documents using LLMs. However, they only investigate how the prompting strategies affect gpt-3.5-turbo and gpt-4-turbo at the inference stage. In our study, we first delve into how these prompting strategies impact the fine-tuning process, as shown in Figure 1, and we present our findings in Section 4.
3.2 Datasets
Parallel Documents
Following Zhang et al. (2022), we conduct experiments on IWSLT2017 translation tasks (Cettolo et al., 2017). IWSLT2017 comprises translation datasets sourced from TED talks, encompassing translations between English and nine other languages, including Arabic, German, French, Italian, Japanese, Korean, Dutch, Romanian, and Chinese. There are approximately $1.9K$ sentence-aligned parallel documents with about $240K$ sentences for each language pair. The dataset statistics can be found in Appendix A.
Monolingual Documents
We gather monolingual documents for all the target languages in our translation tasks, totaling ten languages. To manage computational limitations and address concerns about catastrophic forgetting that might result from excessive continued training, we leverage the data pruning technique suggested by Marion et al. (2023) to select $100M$ tokens for each language, including English, from the CulturaX corpus (Nguyen et al., 2023), totaling $1B$ tokens.
     [<src_lang> Context]: <src1> <src2> <src3>  [<tgt_lang> Context]: <tgt1> <tgt2> <tgt3>  [<src_lang> Sentence]: <src4>  [<tgt_lang> Sentence]: <tgt4>
(a) Prompt 1
     [<src_lang>]: <src1> [<tgt_lang>]: <tgt1>  [<src_lang>]: <src2> [<tgt_lang>]: <tgt2>  [<src_lang>]: <src3> [<tgt_lang>]: <tgt3>  [<src_lang>]: <src4> [<tgt_lang>]: <tgt4>
(b) Prompt 2
     [<src_lang> Context]: <src1> <src2> <src3>  [<tgt_lang> Context]: <tgt1> <tgt2> <tgt3>  Given the provided parallel context, translate the following <src_lang> sentence to <tgt_lang>:  [<src_lang> Sentence]: <src4>  [<tgt_lang> Sentence]: <tgt4>
(c) Prompt 3
     [<src_lang>]: <src1> [<tgt_lang>]: <tgt1>  [<src_lang>]: <src2> [<tgt_lang>]: <tgt2>  [<src_lang>]: <src3> [<tgt_lang>]: <tgt3>  Given the provided parallel sentence pairs, translate the following <src_lang> sentence to <tgt_lang>:  [<src_lang>]: <src4> [<tgt_lang>]: <tgt4>
(d) Prompt 4
Figure 1: Prompt types used in the preliminary study. <src_lang> and <tgt_lang> indicate the source and target languages. <src*> and <tgt*> indicate the source and target sentences. Note that the target sentences <tgt*> are only used during training and are replaced with the hypotheses <hyp*> generated by the model during inference. Concrete examples for each prompt variation can be found in Appendix C.
3.3 Models
Baselines
The baseline models in this study can be classified into three categories, including state-of-the-art LLMs and SenMT models, and our re-implemented DocMT models:
  • State-of-the-art SenMT models: Our selection includes models such as NLLB, which are available with three different sets of parameters: 600M, 1.3B, and 3.3B. [^1] Model signatures: facebook/nllb-200-distilled-600M, facebook/nllb-200-1.3B, and facebook/nllb-200-3.3B. We also incorporate the widely-used commercial translation system, Google Translate.
  • State-of-the-art LLMs: For our baseline LLMs in the context of DocMT, we utilize gpt-3.5-turbo and gpt-4-turbo. [^2] Model signatures: gpt-3.5-turbo-1106 and gpt-4-1106-preview. We use the Prompt 4 as detailed in 1(d) during the translation process.
  • Our re-implemented DocMT models: We conduct full fine-tuning on the concatenation-based DocMT model (Tiedemann and Scherrer, 2017), as well as several recent DocMT baselines (Sun et al., 2022; Wu et al., 2023, 2024b), initialized with mT5 (Xue et al., 2021). These models are available with parameters of 300M, 580M, and 1.2B, representing the strong DocMT baseline.
Ours
In this work, we utilize Llama2-7B, Bloom-7B, and Vicuna-7B, as our backbones. [^3] Llama2 signature: meta-llama/Llama-2-7b-hf, Bloom signature: bigscience/bloom-7b1, and Vicuna signature: lmsys/vicuna-7b-v1.5. Note that Vicuna-v1.5 models are fine-tuned from Llama2. The Llama2 models are predominantly pre-trained on English text, while the Bloom models are pre-trained on multilingual text. The use of Vicuna models allows us to compare the differences between base models and instruction-tuned models (Llama2 vs. Vicuna). We denote those fully fine-tuned models as L-7B-fft, B-7B-fft, and V-7B-fft. We denote those models fine-tuned with LoRA (Hu et al., 2022) as L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA. The optimization details can be found in Appendix B.
3.4 Evaluation
Evaluation Metrics
We evaluate the translation on the IWSLT2017 test sets (Cettolo et al., 2017) using sentence-level BLEU (Papineni et al., 2002) and document-level BLEU (Liu et al., 2020) using SacreBLEU (Post, 2018), denoted as $s$BLEU and $d$BLEU. [^4] BLEU signature: nrefs:1|case:mixed|eff:no|
tok:[13a|ja-mecab-0.996-IPA|ko-mecab-0.996/ko
-0.9.2-KO|zh]|smooth:exp|version:2.3.1
Furthermore, as conventional MT metrics like BLEU demonstrate poor correlation to human judgments (Freitag et al., 2022), we also evaluate the translation quality with the state-of-the-art neural evaluation metric COMET Rei et al. (2020). [^5] COMET signature: Unbabel/wmt22-comet-da. Moreover, we use the average sentence-level BLEU $\mu_{s\textsc{BLEU}}$, the average document-level BLEU $\mu_{d\textsc{BLEU}}$, and the average COMET $\mu_{\textsc{COMET}}$ for the overall performance.
Settings
We evaluate our models in two settings: Bilingual English-from/to-Many Translations and Multilingual English-from/to-Many Translations. For the bilingual translations, we train our models on the training set for one specific language pair. For the multilingual translations, we combine all the available training sets for training.
Inference
In this work, we explore two inference strategies. The first strategy involves using previous translations as context for the current translation, translating test examples in their original order. This begins with the first sentence, which is free from context, as illustrated in 1(d). We refer to this strategy as ReUse. Alternatively, sentences within the context can be re-generated individually for each document translation to prevent error propagation during the translation process. We call this strategy ReGen. For reasons of inference efficiency, ReUse is used as the default inference strategy throughout this study, unless specified otherwise. A comparison of these two inference strategies is provided in Section 6.
PID $\mu_{s\textsc{BLEU}}$ $\mu_{d\textsc{BLEU}}$ $\mu_{\textsc{COMET}}$
L-7B-LoRA 1 15.5 18.2 67.5
2 19.0 21.9 70.7
3 15.8 18.3 69.8
4 20.2 23.4 72.7
B-7B-LoRA 1 19.3 20.5 70.5
2 20.6 23.5 73.6
3 19.8 20.8 73.9
4 23.1 27.3 76.8
V-7B-LoRA 1 19.0 22.4 74.2
2 20.4 23.5 71.6
3 18.3 21.4 70.0
4 22.4 25.7 76.2
Table 1: Overall performance given by L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA on different prompt variations, across four English-centric translation tasks involving German and Chinese. PID indicates the prompt ID in Figure 1. Best results are highlighted in bold.
4 A Preliminary Study on Prompts
The prompt plays a crucial role in LLM research. Recent studies show that an optimal prompt can greatly enhance model performance and reveal unexpected model capabilities (Kojima et al., 2022; Wei et al., 2022b). Hence, our initial focus is on investigating the prompt’s impact during fine-tuning.
Prompt Variations
Displayed in Figure 1, our preliminary study features four prompt types. These designs aim to tackle two research questions: How does context structure impact translation quality? (Prompt 1 vs. Prompt 2) and How do natural language instructions influence translation quality? (Prompt 1 vs. Prompt 3). We also investigate the combined effect of these aspects in Prompt 4.
# of param. # of train. En-X X-En
param. $\mu_{s\textsc{BLEU}}$ $\mu_{d\textsc{BLEU}}$ $\mu_{\textsc{COMET}}$ $\mu_{s\textsc{BLEU}}$ $\mu_{d\textsc{BLEU}}$ $\mu_{\textsc{COMET}}$
State-of-the-art SenMT baselines
NLLB 600M 23.6 27.3 82.3 18.2 22.1 72.8
01.3B 25.7 29.5 83.5 25.0 28.7 78.1
03.3B 26.8 30.5 84.3 25.8 29.4 78.9
GoogleTrans 24.5 28.4 81.6 25.0 28.5 81.2
State-of-the-art LLMs
gpt-3.5-turbo 26.3 30.1 85.3 30.7 34.1 85.5
gpt-4-turbo 27.0 30.7 86.3 31.7 35.1 86.0
LLM backbones
Llama2-7B 7B 02.7 03.5 40.1 04.2 04.4 52.2
Bloom-7B 7B 02.5 02.9 35.5 06.7 07.3 49.4
Vicuna-7B 7B 10.2 12.4 64.7 09.5 09.8 62.9
Re-implemented DocMT baselines
Doc2Doc-mT5 (2017) 300M 300M 17.2 20.2 75.1 19.4 21.2 75.1
580M 580M 18.6 21.5 78.3 20.7 22.5 77.4
01.2B 01.2B 18.4 21.4 79.2 21.5 23.4 78.7
MR-Doc2Sen-mT5 (2022) 01.2B 01.2B 18.8 21.9 79.9 22.0 23.8 79.3
MR-Doc2Doc-mT5 (2022) 01.2B 01.2B 22.5 24.0
DocFlat-mT5 (2023) 01.2B 01.2B 19.2 22.4 80.2 22.2 24.3 79.3
IADA-mT5 (2024c) 01.2B 01.2B 19.3 22.4 80.4 22.1 24.0 79.5
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 00.7B 008M 17.2 20.2 70.8 23.8 25.7 73.7
L-7B-fft 00.7B 00.7B 13.7 16.2 67.4 22.4 24.1 74.0
B-7B-LoRA 00.7B 008M 17.7 20.5 68.5 29.9 33.6 81.4
B-7B-fft 00.7B 00.7B 12.0 13.8 59.6 22.3 24.5 69.9
V-7B-LoRA 00.7B 008M 15.8 18.6 68.8 21.6 23.3 71.4
V-7B-fft 00.7B 00.7B 14.3 16.8 65.0 21.8 23.5 74.3
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 00.7B 008M 13.9 16.9 67.2 17.1 18.4 62.0
L-7B-fft 00.7B 00.7B 17.1 20.4 73.8 18.3 19.5 69.6
B-7B-LoRA 00.7B 008M 10.3 12.3 58.1 23.5 25.7 75.8
B-7B-fft 00.7B 00.7B 16.0 18.6 67.8 27.2 30.5 78.7
V-7B-LoRA 00.7B 008M 12.9 15.7 66.4 17.2 18.6 60.6
V-7B-fft 00.7B 00.7B 13.8 16.8 69.3 18.4 19.4 65.6
Table 2: Overall performance on IWSLT2017. # of param. indicates the number of parameters of the model. # of train. param. indicates the number of trainable parameters of the model. All the LLM approaches use Prompt 4 (1(d)) during inference. Best results are highlighted in bold. Best results in each group are underlined.
Results
Our investigation analyzes prompt variations using three PEFT models (L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA) on four English-centric translation tasks involving German and Chinese. Overall results are presented in Table 1. Comparing Prompt 1 (1(a)) and Prompt 2 (1(b)), we find that models fine-tuned with Prompt 2 generally outperform those with Prompt 1, indicating Prompt 2’s effectiveness in enhancing LLM performance. Regarding our second research question (1(a) vs. 1(c)), we observe varied performance. L-7B-LoRA and B-7B-LoRA perform better with Prompt 3, while V-7B-LoRA performs better with Prompt 1. These results highlight varying impacts of prompt variations across models and suggest natural language instructions are less effective when using instruction-tuned language models as model backbones. Finally, LLMs with Prompt 4 (1(d)) achieve the best overall performance, suggesting a positive compound effect of context structure and instructions.
Conclusion
As expected, the prompt plays a significant role in LLM performance. A well-structured prompt, combining appropriate context and clear instructions, can significantly boost model performance. In this work, we use Prompt 4 (1(d)) in our experiments.
(a) Bilingual English-from-Many translations.
(b) Multilingual English-from-Many translations.
Figure 2: Breakdown results on $s$BLEU, $d$BLEU, and COMET given by L-7B-LoRA, V-7B-LoRA, B-7B-LoRA, Doc2Doc-mT5-1.2B, and gpt-4-turbo for the bilingual (2(a)) and multilingual (2(b)) translation tasks from other languages to English.
5 Main Results
Overall Performance
In our results presented in Table 2, we observe that gpt-4-turbo and gpt-3.5-turbo significantly outperform all other models. Notably, the NLLB variants also demonstrate superior performance among specialized machine translation (MT) models. In the context of DocMT, conventional DocMT models still outperform our LLM-based DocMT models for translations from English to other languages when evaluated using standard MT metrics. Conversely, for translations from other languages to English, our LLM-based DocMT models perform on par or better than conventional DocMT models in $\mu_{s\textsc{BLEU}}$ and $\mu_{d\textsc{BLEU}}$ metrics, while conventional DocMT models maintain superior performance in $\mu_{\textsc{COMET}}$.
LLM-based DocMT Models
As shown in Table 2, models using LoRA generally outperform fully fine-tuned (FFT) LLMs in bilingual translations, while fully fine-tuned LLMs surpass LoRA models in multilingual translations. The LoRA approach, which updates only a limited number of parameters, allows the model to retain the rich linguistic knowledge of the pre-trained LLM while adapting its capabilities to a particular bilingual context. However, in multilingual translation tasks, this parameter-efficient fine-tuning is insufficient to capture the complexities of multiple languages simultaneously. Conversely, fully fine-tuning all the parameters of LLMs enables better capture of multilingual complexity but makes the model prone to overfitting in bilingual translations.
Breakdown Performance
We present the results for translation tasks from other languages to English in Figure 2. For clarity, only the results from our models using LoRA are shown. We observe that LLM-based DocMT models sometimes even surpassing gpt-4-turbo in certain bilingual and multilingual translation tasks, though they fail completely in others. Our manual review indicates that the primary cause of these failures is off-target translation. An in-depth analysis of this issue is provided in Section 6, and a complete breakdown of the results is available in Appendix E.
6 Analysis
In this section, we examine off-target issues with decoding strategy and use GPT-4 to analyze translation errors. We also explore discourse phenomena, training strategies, the data efficiency of parallel documents, and additional evaluations on recent WMT2023 test sets. Additionally, we evaluate the zero-shot crosslingual transfer capabilities of our models and present the results in Appendix I.
$\mu_{\%}$ Ar Ja Ko Zh
L-7B-LoRA 29.2 87.9 25.5 44.2 93.1
L-7B-fft 40.2 87.9 75.5 92.3 93.6
B-7B-LoRA 02.8 02.9 04.0 08.4 01.6
B-7B-fft 28.0 54.1 43.8 70.4 76.4
V-7B-LoRA 32.3 88.2 40.4 35.7 90.5
V-7B-fft 44.7 94.1 98.3 96.6 94.6
Table 3: Off-target rate (%) provided by our LLM-based DocMT models for translation tasks from selective languages to English. $\mu_{\%}$ indicates the average off-target rate across all nine language pairs. A lower off-target rate indicates better performance. The complete results are presented in Appendix F.
Off-Target Translation
In Figure 2, our LLM-based DocMT models excel in some translation tasks but struggle in others due to off-target translation issues. We investigate this problem using the fasttext library (Bojanowski et al., 2017) to identify translation languages and quantify off-target rates, which represent the proportion of translations that are off-target. Results are presented in Table 3, with off-target rates reaching up to 98.3% in failing tasks. Notably, only B-7B-LoRA consistently maintains low off-target rates, likely due to Bloom-7B’s multilingual pre-training. Detailed off-target rates are provided in Appendix F.
$s$BLEU $d$BLEU COMET
L-7B-LoRA + ReUse 03.9 04.1 53.9
L-7B-LoRA + ReGen 17.5 19.6 69.7
L-7B-fft + ReUse 02.5 02.6 51.6
L-7B-fft + ReGen 15.9 17.4 65.0
Table 4: Results on Arabic-English translation given by different decoding strategies.
Inference Strategy
As described in Section 3.4, previous translations provide context for the current translation. We hypothesize that the high off-target translation rate in Table 3 is due to error propagation during the decoding stage. Therefore, we refer to the decoding strategy in Section 3.4 as ReUse, and introduce an alternative strategy where all context translations are re-generated, called ReGen. As shown in Table 4, the ReGen strategy significantly improves translation quality, confirming our hypothesis on decoding error propagation. However, the inference cost of ReGen is four times higher than that of ReUse. These findings highlight the main reason for translation failures in LLM-based DocMT models.
002002002002004004004004006006006006008008008008001,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8002,00020002{,}0002 , 0002,20022002{,}2002 , 2002,40024002{,}4002 , 400MistranslationOvertranslationUndertranslationAdditionOmissionGrammarCohesionCoherenceInconsistent styleMultiple terms in translation1,39513951{,}3951 , 3957177177177171,22412241{,}2241 , 2248408408408404844844844842342342342347097097097094544544544547257257257253463463463461,35613561{,}3561 , 3567177177177171,35813581{,}3581 , 3588638638638636166166166162662662662667277277277275655655655657717717717713393393393391,71217121{,}7121 , 7127157157157159779779779772452452452454174174174174864864864868048048048045925925925928248248248242732732732732,00220022{,}0022 , 0028368368368361,55115511{,}5511 , 551586586586586868868868868474474474474933933933933717717717717854854854854386386386386# of occurencesL-7B-LoRAL-7B-fftGoogleTransDoc2Doc-mT5-1.2B
Figure 3: Error type analysis given by gpt-4-turbo for translations from English to German, Romanian, and Chinese. The error types in orange are context-dependent. We omit those error types that are rare.
Translation Errors
To understand the translation capabilities of our LLM-based DocMT models, we select specific error types from the Multidimensional Quality Metrics (MQM) framework (Burchardt, 2013). Kocmi and Federmann (2023) demonstrate that gpt-4 can identify error spans and achieve state-of-the-art MT evaluation accuracy, so we use gpt-4-turbo to analyze translation errors in texts translated by these models. Due to resource constraints, we focus on four models: L-7B-LoRA, L-7B-fft, Doc2Doc-mT5-1.2B, and GoogleTrans, assessing translations from English to German, Romanian, and Chinese. The error identification prompt is detailed in Appendix D, and we present the frequency of error types in Figure 3. Most errors are limited to individual sentences. Despite similar scores in metrics such as $s$BLEU, $d$BLEU, and COMET among the models, our LLM-based DocMT models (L-7B-LoRA and L-7B-fft) exhibit fewer context-independent and context-dependent errors. This suggests that current evaluation metrics may not sufficiently assess document-level translations and indicates that fine-tuning LLMs holds promise for enhancing DocMT performance.
En-De En-Fr
Doc2Doc-mT5-1.2B 77.0 89.9
L-7B-LoRA 83.1 95.1
L-7B-fft 81.1 94.6
B-7B-LoRA 75.5 91.9
B-7B-fft 68.3 90.8
V-7B-LoRA 84.9 94.8
V-7B-fft 84.4 94.8
Table 5: Accuracy (in %) on the English-German and English-French contrastive test sets. Best results are highlighted in bold.
Discourse Phenomena
To evaluate our LLM-based DocMT model’s ability to leverage contextual information, we assessed it using the English-German contrastive test set by Müller et al. (2018) and the English-French contrastive test set by Lopes et al. (2020). This evaluation tests the model’s accuracy in selecting the correct pronoun from multiple translation options. Results, shown in Table 5, reveal that models initialized with Llama2-7B and Vicuna-7B outperform Doc2Doc-mT5-1.2B, while Bloom-7B-initialized models perform worse, indicating that contextual understanding is mostly acquired during pre-training, as detailed by Scao et al. (2022) due to the lack of German text in Bloom pre-training. The generative accuracy results following Post and Junczys-Dowmunt (2023) are presented in Appendix G.
$s$BLEU $d$BLEU COMET
One-Stage
Nl-En 30.2 32.4 71.2
Ro-En 28.3 30.1 72.3
Ar-En 01.3 02.1 50.1
Zh-En 00.1 00.1 60.9
Two-Stage
Nl-En 38.9 41.9 87.0
Ro-En 38.2 41.4 87.3
Ar-En 02.5 02.6 51.6
Zh-En 00.1 00.1 67.1
Three-Stage
Nl-En 39.1 42.1 87.0
Ro-En 38.4 41.6 87.3
Ar-En 02.3 02.4 52.4
Zh-En 00.3 00.3 67.4
Table 6: Results from one-stage, two-stage, and three-stage training strategies. The one-stage strategy involves directly fine-tuning Llama2-7B on parallel documents. The two-stage results are produced by L-7B-fft. In the three-stage strategy, all model parameters of Llama2-7B are fine-tuned across all three stages.
Training Strategy
In this study, we follow the two-stage approach of Xu et al. (2023). Unlike traditional DocMT methods that typically start with parallel sentence training, we investigate the effectiveness of this strategy on LLM-based DocMT models. Specifically, we attempt to directly fine-tune the LLMs on parallel documents (one-stage) and add an extra fine-tuning stage using parallel sentences to the training strategy in Section 3.1 (three-stage). The results in Table 6 indicate that both the one-stage and three-stage training strategies are sub-optimal for both high-performing languages (Dutch and Romanian) and low-performing languages (Arabic and Chinese) with LLM-based DocMT models.
1102030405060708090100404040406060606080808080Percentage (%) of training dataCOMETL-7B-LoRAL-7B-fftB-7B-LoRAB-7B-fftV-7B-LoRAV-7B-fft
Figure 4: COMET-Percentage (%) of parallel documents for the translations from English to German.
Data Efficiency of Parallel Documents
We explore the data efficiency of parallel documents. Results for English-German translation are presented in Figure 4, and for English-Romanian and English-Chinese in Appendix H. While LLMs typically excel with minimal training data, different fine-tuning strategies show distinct scaling behaviors. Our LoRA models match full training set performance with just $10\%$ of the data (around $20K$ examples), while fully fine-tuned models achieve near-equivalent performance with only about $1\%$ of the data (approximately $2K$ examples). These are crucial for low-resource languages, as recent LLMs are mainly pre-trained on English text.
Evaluation on Recent Test Sets
Given their extensive pre-training on large text corpora, LLMs are susceptible to data leakage risks. We evaluate our models using recent test sets from WMT2023 (Koehn et al., 2023). These tests, conducted between English and German, assess the out-of-domain generalization of our models and help mitigate data leakage risks. We use spaCy to segment documents and discard any parallel documents where the source and target sides have a differing number of sentences. Our findings, presented in Table 7, reveal that while mT5-based models outperform LLM-based models in Table 2, LLM-based models excel in translating out-of-domain text on the WMT2023 test sets. These results highlight the ability of LLM-based DocMT models to generalize well to out-of-domain translation tasks.
En-De De-En
$d$BLEU COMET $d$BLEU COMET
Doc2Doc-mT5-1.2B 20.2 74.4 20.0 76.5
MR-Doc2Sen-mT5 20.5 74.9 21.0 76.5
MR-Doc2Doc-mT5 21.2 75.6 21.5 76.5
DocFlat-mT5 20.9 75.1 21.8 76.5
IADA-mT5 21.2 75.4 22.0 76.5
L-7B-LoRA 28.9 76.4 35.5 83.2
L-7B-fft 29.0 77.0 36.1 84.0
B-7B-LoRA 23.7 73.0 30.5 80.8
B-7B-fft 21.0 69.0 30.0 80.5
V-7B-LoRA 20.5 63.8 33.9 81.8
V-7B-fft 27.8 75.0 34.7 83.1
Table 7: $d$BLEU and COMET on WMT2023 test sets. Best results are highlighted in bold.
7 Conclusion
In this study, we adapt large language models (LLMs) for document-level machine translation (DocMT) across specific language pairs and found that moderately-sized, fine-tuned models can sometimes surpass larger models like GPT-4 in translation performance. By investigating prompt strategies and conducting extensive experiments with different fine-tuning methods and LLM backbones, we highlight both the potential and challenges of using LLMs for DocMT. Despite achieving promising results, issues like off-target translations due to error propagation persist. Our analysis of translation errors, discourse phenomena, strategies for training and inference, and data efficiency underscores the strengths and limitations of LLM-based DocMT models, providing a foundation for future research to enhance their effectiveness.
8 Limitations
Constraints on Model Scale
Our research is confined to language models of a moderate size, specifically those with $7B$ parameters. This limitation is due to the constraints of our available resources. Consequently, it is crucial to acknowledge that the outcomes of our study might vary if conducted with larger models.
Instability in Training
The process of supervised fine-tuning for LLMs shows instability in our observations. As detailed in Figure 4, there are noticeable inconsistencies in performance. These variations are too significant to attribute solely to the randomness inherent in training. In some cases, the fine-tuning of LLMs fails to reach convergence. Unfortunately, our limited resources restrict us from investigating these failures in depth or devising potential remedies.
Influence of Prompting Techniques
Section 4 of our study highlights the significant role of prompting methods in fine-tuning. We experiment with four different prompting techniques. It is important to note that the prompt we recommend may not be the most effective, potentially leading to suboptimal performance of our models.
We acknowledge these limitations and leave them to the future work.
Appendix A Statistics of Parallel Documents
We present the dataset statistics of parallel documents in Table 8.
Appendix B Optimization and Hyperparameters
Fine-tuning on Monolingual Documents
We fine-tune all the parameters of large language models (LLMs) using a learning rate of $5\times 10^{-5}$ and a batch size of $256$. During the training process, we apply the linear learning rate schedule, which includes a warm-up phase comprising $10\%$ of the total training steps.
Fine-tuning on Parallel Documents
When fine-tuning L-7B-LoRA and V-7B-LoRA on parallel documents, we use a learning rate of $5\times 10^{-5}$ and a batch size of $64$. We apply a linear learning rate schedule with a 10% warm-up phase. The LoRA rank is set to 16, affecting only 0.1% of the parameters (about 8M parameters). The same hyperparameters are used for fine-tuning Doc2Doc-mT5 models, except for a learning rate of $5\times 10^{-4}$. L-7B-LoRA and V-7B-LoRA are fine-tuned for up to $3$ epochs, and Doc2Doc-mT5 models for up to $10$ epochs. Early stopping is based on validation loss.
     [English Context]: And it’s truly a great honor to have the opportunity to come to this stage twice; I’m extremely grateful. I have been blown away by this conference, and I want to thank all of you for the many nice comments about what I had to say the other night. And I say that sincerely, partly because  I need that.  [German Context]: Es ist mir wirklich eine Ehre, zweimal auf dieser Bühne stehen zu dürfen. Tausend Dank dafür. Ich bin wirklich begeistert von dieser Konferenz, und ich danke Ihnen allen für die vielen netten Kommentare zu meiner Rede vorgestern Abend. Das meine ich ernst, teilweise deshalb -- weil ich es wirklich brauchen kann!  [English Sentence]: Put yourselves in my position.  [German Sentence]: Versetzen Sie sich mal in meine Lage!
(a) Prompt 1
     [English]: And it’s truly a great honor to have the opportunity to come to this stage twice; I’m extremely grateful.  [German]: Es ist mir wirklich eine Ehre, zweimal auf dieser Bühne stehen zu dürfen. Tausend Dank dafür.  [English]: I have been blown away by this conference, and I want to thank all of you for the many nice comments about what I had to say the other night.  [German]: Ich bin wirklich begeistert von dieser Konferenz, und ich danke Ihnen allen für die vielen netten Kommentare zu meiner Rede vorgestern Abend.  [English]: And I say that sincerely, partly because  I need that.  [German]: Das meine ich ernst, teilweise deshalb -- weil ich es wirklich brauchen kann!  [English]: Put yourselves in my position.  [German]: Versetzen Sie sich mal in meine Lage!
(b) Prompt 2
     [English Context]: And it’s truly a great honor to have the opportunity to come to this stage twice; I’m extremely grateful. I have been blown away by this conference, and I want to thank all of you for the many nice comments about what I had to say the other night. And I say that sincerely, partly because  I need that.  [German Context]: Es ist mir wirklich eine Ehre, zweimal auf dieser Bühne stehen zu dürfen. Tausend Dank dafür. Ich bin wirklich begeistert von dieser Konferenz, und ich danke Ihnen allen für die vielen netten Kommentare zu meiner Rede vorgestern Abend. Das meine ich ernst, teilweise deshalb -- weil ich es wirklich brauchen kann!  Given the provided parallel context, translate the following English sentence to German:  [English Sentence]: Put yourselves in my position.  [German Sentence]: Versetzen Sie sich mal in meine Lage!
(c) Prompt 3
     [English]: And it’s truly a great honor to have the opportunity to come to this stage twice; I’m extremely grateful.  [German]: Es ist mir wirklich eine Ehre, zweimal auf dieser Bühne stehen zu dürfen. Tausend Dank dafür.  [English]: I have been blown away by this conference, and I want to thank all of you for the many nice comments about what I had to say the other night.  [German]: Ich bin wirklich begeistert von dieser Konferenz, und ich danke Ihnen allen für die vielen netten Kommentare zu meiner Rede vorgestern Abend.  [English]: And I say that sincerely, partly because  I need that.  [German]: Das meine ich ernst, teilweise deshalb -- weil ich es wirklich brauchen kann!  Given the provided parallel sentence pairs, translate the following English sentence to German:  [English]: Put yourselves in my position.  [German]: Versetzen Sie sich mal in meine Lage!
(d) Prompt 4
Figure 5: Prompt types used in the preliminary study. <src_lang> and <tgt_lang> indicate the language IDs. <src*> and <tgt*> indicate the source and target sentences. Note that the target sentences <tgt*> are only used during training and are replaced with the hypotheses <hyp*> generated by the model during inference.
Appendix C Prompt Types
We present concrete examples of prompt variations in Figure 5.
     [Context]:    [Source]: <src1>  [Reference]: <tgt1>  [Hypothesis]: <hyp1>  [Source]: <src2>  [Reference]: <tgt2>  [Hypothesis]: <hyp2>  [Source]: <src3>  [Reference]: <tgt3>  [Hypothesis]: <hyp3>    [Current Sentence]:    [Source]: <src4>  [Reference]: <tgt4>  [Hypothesis]: <hyp4>    [Error Types]:    - Mistranslation: Error occuring when the target content does not accurately represent the source content.  - Overtranslation: Error occuring in the target content that is inappropriately more specific than the source content.  - Undertranslation: Error occuring in the target content that is inappropriately less specific than the source content.  - Addition: Error occuring in the target content that includes content not present in the source.  - Omission: Error where content present in the source is missing in the target.  - Unjustfied euphemism: Target content that is potentially offensive in some way in the source language, but that has been inappropriately "watered down" in the translation.  - Do not translate: Error occuring when a text segment marked "Do not translate!" is translated in the target text.  - Untranslated: Error occuring when a text segment that was intended for translation is omitted in the target content.  - Retained factual error: Untrue statement or an incorrect data value present in the source content and retained in the target content.  - Completeness: Source text incomplete, resulting in instances where needed content is missing in the source language.  - Grammar: Error that occurs when a text string (sentence, phrase, other) in the translation violates the grammatical rules of the target language.  - Punctuation: Punctuation incorrect according to target language conventions.  - Spelling: Error occurring when a word is misspelled.  - Duplication: Content (e.g., a word or longer portion of text) repeated unintentionally.  - Unclear reference: Relative pronouns or other referential mechanisms unclear in their reference.  - Cohesion: Portions of the text needed to connect it into an understandable whole (e.g., reference, substitution, ellipsis, conjunction, and lexical cohesion) missing or incorrect.  - Coherence: Text lacking a clear semantic relationship between its parts, i.e., the different parts don’t hang together, don’t follow the discourse conventions of the target language, or don’t "make sense."  - Inconsistent style: Style that varies inconsistently throughout the text, e.g., One part of a text is written in a clear, "terse" style, while other sections are written in a more wordy style.  - Multiple terms in translation: Error where source content terminology is correct, but target content terms are not used consistently.    Considering the provided context, please identify the errors of the translation from the source to the target in the current sentence based on a subset of Multidimensional Quality Metrics (MQM) error typology.  You should pay extra attention to the error types related to the relationship between the current sentence and its context, such as "Unclear reference", "Cohesion", "Coherence", "Inconsistent style", and "Multiple terms in translation".  You should list all the errors you find in the sentence, and provide a justification for each error.  Your output should always be in JSON format, formatted as follows: {’justification’: ’...’, ’error_types’: [...]}.
Figure 6: Prompt used for analyzing translation error types.
Appendix D GPT-4 Prompts
We present the prompts used for error type analysis in Figure 6.
$\mu_{s\textsc{BLEU}}$ Ar De Fr It Ja Ko Nl Ro Zh
State-of-the-art SenMT baselines
NLLB-600M 23.6 14.2 22.2 38.5 36.0 11.4 16.2 30.2 25.6 17.6
NLLB-1.3B 25.7 16.2 27.6 40.6 37.7 12.6 18.6 32.2 27.3 18.3
NLLB-3.3B 26.8 17.4 28.8 41.3 39.2 14.1 19.5 33.7 28.1 18.7
GoogleTrans 24.5 14.2 25.3 38.0 35.0 11.6 16.5 29.6 24.0 26.4
State-of-the-art LLMs
gpt-3.5-turbo 26.3 14.9 27.2 40.5 36.6 13.2 15.9 31.5 26.6 30.4
gpt-4-turbo 27.0 16.1 27.4 40.0 35.8 14.1 18.3 32.2 27.3 31.6
LLM backbones
Llama2 02.8 00.4 06.6 04.5 01.0 01.0 01.9 00.2 01.9 07.4
Bloom 02.5 01.0 01.0 12.1 01.4 00.1 03.1 00.7 00.1 03.4
Vicuna 10.2 04.5 06.4 06.4 08.6 10.2 09.8 13.9 06.8 25.4
Re-implemented DocMT baselines
Doc2Doc-mT5-300M 17.2 09.4 16.8 24.0 21.0 11.0 13.7 20.5 17.1 21.6
Doc2Doc-mT5-580M 18.6 10.8 18.2 24.9 23.0 12.9 15.2 21.7 17.8 22.9
Doc2Doc-mT5-1.2B 18.4 10.3 18.1 24.9 22.4 13.9 15.4 19.6 18.8 22.6
MR-Doc2Sen-mT5 18.8 10.2 18.8 25.6 22.3 14.5 16.2 19.6 19.3 22.8
MR-Doc2Doc-mT5
DocFlat-mT5 19.2 11.0 19.2 25.7 22.6 14.7 16.5 20.3 19.2 23.8
IADA-mT5 19.7 11.7 19.4 26.3 23.9 15.2 16.9 20.9 19.6 23.4
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 17.2 13.0 25.1 34.9 06.8 08.7 13.0 03.7 22.7 27.3
L-7B-fft 13.7 13.1 25.3 19.5 02.6 07.9 07.2 04.1 21.1 22.8
B-7B-LoRA 17.7 12.1 20.6 32.6 32.9 03.6 01.4 28.1 12.2 15.7
B-7B-fft 12.0 10.1 19.6 38.5 00.1 01.9 02.4 01.5 19.9 14.5
V-7B-LoRA 16.4 13.3 20.1 20.7 13.6 09.1 14.3 05.5 23.0 28.1
V-7B-fft 14.3 13.5 23.3 21.1 04.8 03.8 15.9 03.3 17.4 25.8
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 13.9 08.8 22.2 27.9 05.0 07.7 09.4 03.4 19.4 21.6
L-7B-fft 17.1 07.8 22.0 32.5 17.9 12.5 11.5 04.2 20.3 25.1
B-7B-LoRA 10.3 12.5 03.8 36.0 12.8 05.8 01.6 03.7 03.4 12.9
B-7B-fft 16.0 12.6 15.8 37.4 27.7 02.0 01.3 18.4 07.8 21.3
V-7B-LoRA 12.9 08.0 15.4 28.2 06.2 10.1 07.1 03.3 15.1 22.4
V-7B-fft 13.8 10.4 15.2 11.3 08.4 11.1 12.2 12.4 17.0 26.3
Table 9: Breakdown $s$BLEU results for the translation tasks from English to other languages.
$\mu_{d\textsc{BLEU}}$ Ar De Fr It Ja Ko Nl Ro Zh
State-of-the-art SenMT baselines
NLLB-600M 27.3 15.4 26.0 42.1 39.0 17.1 23.8 33.8 28.5 20.2
NLLB-1.3B 29.5 17.4 31.8 44.0 40.8 18.1 26.6 35.8 30.2 21.0
NLLB-3.3B 30.5 18.6 32.9 44.6 42.2 19.8 27.4 37.0 30.9 21.2
GoogleTrans 28.4 16.0 29.3 41.3 38.5 15.7 23.4 32.8 26.7 32.1
State-of-the-art LLMs
gpt-3.5-turbo 30.1 16.4 30.9 43.7 39.7 17.5 22.7 34.5 29.0 36.3
gpt-4-turbo 30.7 17.4 31.1 43.2 39.0 18.4 25.3 35.3 29.8 37.2
LLM backbones
Llama2 03.5 00.5 07.4 04.9 01.1 01.8 03.7 00.2 02.2 09.6
Bloom 02.8 01.0 01.3 12.9 01.7 00.3 02.7 01.0 00.1 04.4
Vicuna 12.4 05.7 06.4 07.1 08.5 15.0 16.1 14.4 07.4 31.2
Re-implemented DocMT baselines
Doc2Doc-mT5-300M 20.2 10.3 18.8 26.1 21.9 16.8 21.5 21.5 18.4 26.6
Doc2Doc-mT5-580M 21.5 11.7 20.0 27.0 23.9 18.4 23.2 22.6 18.8 28.0
Doc2Doc-mT5-1.2B 21.4 11.2 20.1 27.1 23.4 19.7 23.0 20.7 20.2 27.1
MR-Doc2Sen-mT5
MR-Doc2Doc-mT5 21.9 11.9 20.7 27.9 23.8 19.8 23.3 21.5 20.7 27.9
DocFlat-mT5 22.5 12.1 20.8 28.0 24.7 20.9 24.3 21.9 21.7 27.9
IADA-mT5 22.4 12.2 21.3 28.4 24.0 20.8 24.1 21.1 21.4 28.1
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 20.2 14.7 29.1 37.5 07.3 13.9 19.5 04.2 22.9 33.1
L-7B-fft 16.2 14.7 29.4 20.6 02.7 12.5 12.3 04.5 21.6 27.6
B-7B-LoRA 20.5 13.7 24.8 36.1 36.3 06.9 02.6 32.2 12.2 19.7
B-7B-fft 13.8 11.2 23.6 41.7 00.1 03.7 03.9 01.7 20.1 18.1
V-7B-LoRA 19.3 14.9 23.1 21.8 14.7 14.3 21.6 05.9 23.3 34.2
V-7B-fft 16.8 15.2 26.9 22.3 04.9 06.2 23.6 03.7 17.5 31.1
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 16.9 10.2 26.0 30.4 05.6 12.8 15.7 03.8 20.2 27.6
L-7B-fft 20.4 09.2 25.5 35.5 18.5 19.2 18.3 04.9 20.9 31.6
B-7B-LoRA 12.3 14.3 04.4 39.5 14.4 10.8 03.3 03.9 03.5 16.8
B-7B-fft 18.6 14.2 19.5 40.9 30.9 03.8 02.6 20.9 08.5 26.5
V-7B-LoRA 15.7 09.5 18.1 31.0 06.5 16.0 11.8 03.9 16.1 28.1
V-7B-fft 16.8 12.0 17.5 12.9 08.8 17.4 19.5 12.4 17.8 32.7
Table 10: Breakdown $d$BLEU results for the translation tasks from English to other languages.
$\mu_{\textsc{COMET}}$ Ar De Fr It Ja Ko Nl Ro Zh
State-of-the-art SenMT baselines
NLLB-600M 82.3 82.6 81.3 83.7 86.3 79.5 82.5 84.0 85.0 75.4
NLLB-1.3B 83.5 84.3 83.0 84.8 87.5 80.1 84.2 85.2 86.3 76.4
NLLB-3.3B 84.3 84.8 84.1 85.3 88.0 82.1 85.2 86.0 86.4 76.7
GoogleTrans 81.6 81.5 80.2 82.3 85.1 80.7 79.9 83.7 82.9 78.5
State-of-the-art LLMs
gpt-3.5-turbo 85.3 83.8 84.6 85.9 87.7 84.7 84.2 86.3 86.5 83.9
gpt-4-turbo 86.3 85.4 85.4 86.2 87.8 86.0 86.4 87.0 87.4 84.8
LLM backbones
Llama2 40.1 37.4 41.5 41.2 39.3 39.7 42.8 35.5 41.6 42.1
Bloom 35.5 34.9 34.8 45.3 33.0 33.9 35.5 34.2 28.6 39.0
Vicuna 64.7 68.3 48.7 49.0 62.5 81.3 72.4 64.1 56.1 80.2
Re-implemented DocMT baselines
Doc2Doc-mT5-300M 75.1 77.2 71.0 72.4 74.1 78.5 77.8 74.3 74.0 77.0
Doc2Doc-mT5-580M 78.3 80.8 74.4 74.8 77.9 81.5 82.0 76.8 76.6 79.6
Doc2Doc-mT5-1.2B 79.2 81.1 75.9 75.9 78.9 82.9 82.4 76.5 79.3 80.3
MR-Doc2Sen-mT5 79.9 82.2 76.3 76.5 80.0 83.6 83.1 76.7 79.7 80.9
MR-Doc2Doc-mT5
DocFlat-mT5 80.4 81.8 77.3 76.9 80.3 83.6 83.4 78.0 80.7 81.8
IADA-mT5 80.7 82.4 77.0 77.3 80.9 84.1 83.7 77.8 80.9 81.8
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 70.8 82.5 80.3 79.1 42.9 70.4 75.4 42.5 83.6 80.6
L-7B-fft 67.4 83.1 82.0 59.4 38.8 65.8 69.7 49.2 82.7 75.5
B-7B-LoRA 68.5 77.0 75.4 76.8 85.1 51.4 40.4 82.9 61.7 65.5
B-7B-fft 59.6 68.4 72.6 83.8 45.3 40.3 45.2 46.8 71.6 62.2
V-7B-LoRA 69.7 82.7 70.8 60.1 56.2 70.2 76.9 44.1 85.0 81.3
V-7B-fft 65.0 83.1 73.9 58.8 41.5 54.8 81.1 42.4 69.6 79.4
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 67.2 78.8 74.1 69.3 40.9 65.9 74.7 44.9 79.3 77.1
L-7B-fft 73.8 78.8 76.3 77.0 65.2 80.6 79.0 45.3 81.9 79.8
B-7B-LoRA 58.1 80.9 43.8 79.9 54.9 66.7 46.6 38.1 46.9 65.3
B-7B-fft 67.8 82.2 69.6 83.8 81.7 46.3 45.7 70.5 54.5 76.0
V-7B-LoRA 66.4 78.2 65.5 70.8 46.9 74.5 66.4 41.9 76.0 77.6
V-7B-fft 69.3 80.3 64.7 54.2 52.0 78.1 77.5 57.6 78.1 81.1
Table 11: Breakdown COMET results for the translation tasks from English to other languages.
$\mu_{s\textsc{BLEU}}$ Ar De Fr It Ja Ko Nl Ro Zh
State-of-the-art SenMT baselines
NLLB-600M 18.2 26.8 11.0 31.0 22.7 10.9 13.6 17.1 13.2 17.9
NLLB-1.3B 25.0 35.9 18.8 37.4 35.3 13.2 15.7 26.0 23.1 19.9
NLLB-3.3B 25.8 36.5 22.3 36.8 33.5 12.4 18.5 28.3 25.2 19.1
GoogleTrans 25.0 28.7 26.1 34.7 35.1 10.2 13.3 30.8 29.6 16.6
State-of-the-art LLMs
gpt-3.5-turbo 30.7 35.8 30.8 40.7 41.8 15.5 17.3 36.1 35.4 22.9
gpt-4-turbo 31.7 37.2 31.2 41.4 42.3 15.9 19.8 36.6 36.5 24.4
LLM backbones
Llama2 04.2 00.1 06.2 01.3 04.4 00.0 00.1 15.1 10.3 00.1
Bloom 06.7 04.7 08.7 14.3 16.9 00.1 00.3 09.4 05.5 00.2
Vicuna 09.5 01.1 14.4 24.9 14.1 04.7 00.4 17.2 08.5 00.0
Re-implemented DocMT baselines
Doc2Doc-mT5-300M 19.4 23.0 19.5 26.6 25.4 09.5 11.4 22.0 22.6 14.5
Doc2Doc-mT5-580M 20.7 24.2 20.3 28.0 26.6 11.0 11.6 24.4 23.8 16.1
Doc2Doc-mT5-1.2B 21.5 25.7 21.0 28.7 27.3 11.0 12.8 25.4 25.0 16.8
MR-Doc2Sen-mT5 22.5 26.9 22.0 29.9 27.7 11.8 13.9 26.5 26.1 18.0
MR-Doc2Doc-mT5
DocFlat-mT5 22.5 26.6 22.3 29.7 28.4 11.7 13.9 26.4 25.9 17.6
IADA-mT5 23.1 26.9 22.7 30.5 28.5 12.8 14.5 27.1 26.1 18.7
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 23.8 03.9 33.1 40.3 45.2 08.3 05.0 39.2 39.0 00.1
L-7B-fft 22.4 02.5 32.2 42.6 44.8 01.0 01.0 38.9 38.2 00.1
B-7B-LoRA 29.9 30.9 30.5 41.1 41.2 13.9 15.5 35.0 35.2 25.6
B-7B-fft 22.3 17.0 29.8 41.3 40.7 00.4 01.1 35.6 34.3 01.0
V-7B-LoRA 21.6 03.8 30.8 43.1 29.4 05.5 04.0 39.1 38.7 00.3
V-7B-fft 24.3 24.2 31.0 43.4 45.0 00.0 00.7 36.2 38.0 00.1
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 17.1 02.3 15.7 18.6 41.6 00.4 00.8 37.6 36.7 00.3
L-7B-fft 18.3 15.8 10.3 22.0 40.7 02.1 06.0 29.7 30.8 07.8
B-7B-LoRA 23.5 34.4 28.5 39.5 40.0 00.1 00.6 33.7 34.1 00.6
B-7B-fft 27.2 35.9 28.3 39.7 38.8 07.1 11.2 32.4 33.5 17.9
V-7B-LoRA 17.2 06.1 05.5 12.2 43.3 02.9 01.8 35.9 36.4 10.2
V-7B-fft 18.4 13.0 17.7 16.7 37.6 02.5 03.7 36.2 35.2 03.0
Table 12: Breakdown $s$BLEU results for the translation tasks from other languages to English.
$\mu_{d\textsc{BLEU}}$ Ar De Fr It Ja Ko Nl Ro Zh
State-of-the-art SenMT baselines
NLLB-600M 22.0 30.5 14.7 34.1 26.3 14.8 18.1 21.0 16.8 22.2
NLLB-1.3B 28.6 39.2 22.6 40.1 38.7 17.1 20.3 29.4 26.7 23.7
NLLB-3.3B 29.4 39.7 26.1 39.6 37.0 16.5 23.2 31.3 28.8 22.8
GoogleTrans 28.5 32.0 29.8 37.8 38.9 13.3 17.7 33.7 33.1 20.4
State-of-the-art LLMs
gpt-3.5-turbo 34.0 38.8 34.0 43.4 44.8 19.5 22.0 38.8 38.4 26.9
gpt-4-turbo 35.1 40.2 34.5 44.2 46.3 19.7 24.4 39.3 39.6 28.1
LLM backbones
Llama2 04.4 00.1 06.9 01.5 04.7 00.0 00.1 15.7 10.9 00.1
Bloom 07.3 05.3 09.8 15.2 17.6 00.1 00.5 10.6 06.6 00.2
Vicuna 09.8 01.1 14.5 24.6 14.4 05.9 00.5 18.5 08.9 00.0
Re-implemented DocMT baselines
Doc2Doc-mT5-300M 21.2 24.5 21.1 27.5 26.5 12.6 14.1 23.5 23.9 17.0
Doc2Doc-mT5-580M 22.5 25.5 22.1 28.9 27.8 14.0 14.6 25.8 25.2 18.5
Doc2Doc-mT5-1.2B 23.4 26.9 23.0 29.7 28.4 14.2 15.7 26.8 26.3 19.4
MR-Doc2Sen-mT5 24.0 27.4 24.2 30.3 29.4 14.9 16.1 27.5 26.8 19.8
MR-Doc2Doc-mT5 24.6 28.3 24.3 30.5 29.8 15.7 16.8 27.8 27.8 20.8
DocFlat-mT5 24.6 27.6 24.5 31.1 29.7 15.1 17.0 28.1 27.8 20.3
IADA-mT5 24.5 28.2 24.6 30.9 29.6 15.0 17.1 27.8 27.1 20.5
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 25.7 04.1 36.2 42.2 48.5 10.0 05.9 42.2 42.1 00.1
L-7B-fft 24.1 02.6 35.3 45.1 48.1 01.0 01.1 41.9 41.4 00.1
B-7B-LoRA 33.6 33.0 34.2 44.0 44.9 18.8 20.4 38.1 38.6 30.4
B-7B-fft 24.5 18.6 33.4 44.2 44.4 00.6 01.4 38.7 37.9 01.1
V-7B-LoRA 23.3 03.8 33.9 45.5 30.6 06.7 04.8 42.1 42.0 00.3
V-7B-fft 23.5 02.3 34.1 46.0 48.3 00.0 00.7 39.2 41.0 00.0
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 18.4 02.2 17.6 19.5 44.6 00.4 00.8 40.4 39.7 00.2
L-7B-fft 19.5 15.4 11.1 23.3 43.5 02.4 07.0 31.4 33.0 08.5
B-7B-LoRA 25.7 37.9 32.1 42.5 43.7 00.2 00.7 36.7 37.5 00.5
B-7B-fft 30.5 39.5 31.9 42.6 41.9 10.0 16.0 35.3 37.0 20.6
V-7B-LoRA 18.6 06.4 06.0 13.5 46.6 03.5 02.1 38.3 39.7 11.6
V-7B-fft 19.4 13.4 17.1 17.4 39.6 02.9 04.4 37.8 38.1 03.6
Table 13: Breakdown $d$BLEU results for the translation tasks from other languages to English.
$\mu_{\textsc{COMET}}$ Ar De Fr It Ja Ko Nl Ro Zh
State-of-the-art SenMT baselines
NLLB-600M 72.8 76.6 63.0 79.8 72.4 74.2 76.3 68.4 67.0 77.6
NLLB-1.3B 78.1 82.3 71.9 83.8 80.6 75.3 78.4 77.5 75.4 77.9
NLLB-3.3B 78.9 82.6 76.1 83.7 80.2 74.9 80.1 78.7 76.5 77.7
GoogleTrans 81.2 81.1 82.3 84.6 84.5 75.6 76.9 84.2 83.8 78.1
State-of-the-art LLMs
gpt-3.5-turbo 85.5 85.7 86.0 87.9 88.1 81.4 82.4 87.2 87.4 83.6
gpt-4-turbo 86.0 86.5 86.3 88.2 88.5 81.9 83.5 87.5 87.9 84.2
LLM backbones
Llama2 52.2 50.3 47.1 42.6 51.6 56.1 55.7 56.5 53.0 56.9
Bloom 49.4 50.6 52.8 55.5 56.8 44.5 44.3 51.2 45.2 43.3
Vicuna 62.7 51.3 65.2 70.5 57.3 69.5 56.3 68.0 58.8 67.5
Re-implemented DocMT baselines
Doc2Doc-mT5-300M 75.1 75.0 75.2 78.0 77.5 71.8 72.4 75.3 76.6 73.9
Doc2Doc-mT5-580M 77.4 77.4 77.0 79.7 79.8 74.7 74.3 78.5 79.1 75.8
Doc2Doc-mT5-1.2B 78.7 79.0 78.8 80.9 80.5 75.5 75.8 80.3 80.5 76.8
MR-Doc2Sen-mT5 79.8 80.3 79.8 82.3 81.1 76.4 76.6 81.5 81.8 78.2
MR-Doc2Doc-mT5
DocFlat-mT5 80.3 80.2 80.0 82.5 81.7 77.4 77.6 82.2 82.3 78.4
IADA-mT5 80.4 80.3 80.7 82.9 82.3 77.7 77.7 81.8 81.7 78.5
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 73.7 53.9 84.0 84.1 88.2 59.0 53.0 87.0 87.6 66.9
L-7B-fft 74.0 51.6 81.9 86.5 88.3 63.6 52.9 87.0 87.3 67.1
B-7B-LoRA 81.4 73.3 83.6 87.0 87.1 74.0 73.8 84.8 86.0 82.6
B-7B-fft 69.9 53.7 83.2 86.9 86.8 50.5 43.3 85.1 84.4 55.5
V-7B-LoRA 71.4 54.5 82.6 87.2 64.9 57.8 53.7 87.1 87.3 67.7
V-7B-fft 77.0 77.0 81.3 87.0 88.2 65.6 55.5 84.2 87.2 67.0
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 62.0 58.9 54.0 63.8 83.7 42.5 54.5 73.7 77.2 50.0
L-7B-fft 69.6 52.9 63.2 63.2 85.1 65.3 57.0 86.3 86.3 66.8
B-7B-LoRA 78.7 84.0 82.5 86.5 84.6 62.4 68.4 82.3 84.2 73.8
B-7B-fft 75.8 81.0 83.3 86.2 86.9 59.1 51.7 83.9 85.7 64.4
V-7B-LoRA 60.6 52.4 58.2 57.3 78.8 43.6 47.6 83.3 83.8 40.6
V-7B-fft 65.6 50.9 47.4 53.6 87.2 61.4 56.8 83.3 86.3 63.5
Table 14: Breakdown COMET results for the translation tasks from other languages to English.
Appendix E Breakdown Results
We provide detailed breakdowns of the translation tasks from English to other languages, evaluated using $s$BLEU, $d$BLEU, and COMET. These are presented in Table 9, Table 10, and Table 11, respectively. Additionally, we present similar breakdowns for translations from other languages to English, assessed using the same metrics. These results can be found in Table 12, Table 13, and Table 14.
$\mu_{\%}$ Ar De Fr It Ja Ko Nl Ro Zh
L-7B-LoRA 06.2 00.4 04.6 01.2 00.5 17.1 00.4 00.8 01.0 29.7
L-7B-fft 09.4 00.3 00.9 00.3 04.4 17.7 09.4 15.4 01.2 34.7
B-7B-LoRA 11.2 08.4 01.0 20.8 03.9 16.4 00.0 02.8 00.9 46.9
B-7B-fft 31.8 36.6 15.8 02.7 90.1 10.7 00.1 82.0 00.2 47.7
V-7B-LoRA 10.6 00.2 15.4 00.3 13.3 15.9 00.5 20.3 00.9 28.9
V-7B-fft 08.9 00.1 14.4 00.5 00.4 27.8 00.5 04.6 00.4 31.5
Table 15: Off-target rate (%) provided by our LLM-based DocMT models for translation tasks from English to other languages. $\mu_{\%}$ indicates the average off-target rate. A lower off-target rate indicates better performance.
$\mu_{\%}$ Ar De Fr It Ja Ko Nl Ro Zh
L-7B-LoRA 29.2 87.9 02.0 04.9 01.4 25.5 44.2 01.9 01.8 93.1
L-7B-fft 40.2 87.9 05.8 02.1 01.5 75.5 92.3 01.6 01.9 93.6
B-7B-LoRA 02.8 02.9 02.1 01.0 01.3 04.0 08.4 01.9 02.0 01.6
B-7B-fft 28.0 54.1 02.0 01.0 01.1 43.8 70.4 01.6 01.9 76.4
V-7B-LoRA 32.3 88.2 02.6 01.2 28.0 40.4 35.7 01.9 01.9 90.5
V-7B-fft 44.7 94.1 09.0 01.3 01.3 98.3 96.6 05.3 01.9 94.6
Table 16: Off-target rate (%) provided by our LLM-based DocMT models for translation tasks from other languages to English. $\mu_{\%}$ indicates the average off-target rate. A lower off-target rate indicates better performance.
Appendix F Off-Target Translation
We present the complete results on the off-target translation problem in Table 15 and Table 16.
En-De En-Fr
Doc2Doc-mT5-1.2B 39.8 26.2
L-7B-LoRA 64.4 29.9
L-7B-fft 62.4 28.4
B-7B-LoRA 58.3 23.6
B-7B-fft 49.8 25.1
V-7B-LoRA 48.9 30.2
V-7B-fft 43.5 27.2
Table 17: Generative accuracy (in %) on the English-German and English-French contrastive test sets. Best results are highlighted in bold.
Appendix G Discourse Phenomena
Post and Junczys-Dowmunt (2023) propose to evaluate the accuracy on the constrastive test sets in a generative way. Hence, we present the generative accuracy in Table 17.
1204060801000020202020404040406060606080808080100100100100Percentage (%) of training dataCOMETEnglish-RomanianL-7B-LoRAL-7B-fftB-7B-LoRAB-7B-fftV-7B-LoRAV-7B-fft
1204060801000020202020404040406060606080808080100100100100Percentage (%) of training dataCOMETEnglish-ChineseL-7B-LoRAL-7B-fftB-7B-LoRAB-7B-fftV-7B-LoRAV-7B-fft
Figure 7: COMET-Percentage (%) of training data for the translations from English to Romanian, and Chinese.
Appendix H Scaling Law of Parallel Documents from English to Romanian and Chinese
In Section 6, we find that our LLM-based DocMT models are highly efficient in terms of the amount of training data. To confirm our findings in Section 6, we conduct additional experiments on the translation tasks from English to Romanian and Chinese. As shown in Figure 7, we can confirm the superiority of LLM-based DocMT models with regard to data efficiency.
Appendix I Zero-Shot Crosslingual Transfer
$\mu_{\Delta}$ Ar De Fr It Ja Ko Nl Ro Zh
L-7B-LoRA +29.4 +36.3 +38.8 +37.2 +32.1 +15.9 +17.1 +21.7 +35.8 +29.5
L-7B-fft +29.0 +41.2 +40.5 +37.1 +18.0 +27.7 +29.4 +11.2 +18.5 +37.5
B-7B-LoRA +20.3 0+7.5 +40.7 +20.7 +21.9 +17.5 +15.9 +23.7 +25.3 0+9.8
B-7B-fft +27.3 +14.8 +37.8 +28.9 +43.3 +13.1 +15.3 +38.5 +34.7 +19.5
V-7B-LoRA 0.-8.9 .-12.6 +22.1 +18.9 .-28.6 .-27.8 .-18.7 .-11.8 +12.1 .-34.1
V-7B-fft 0.-1.4 0+7.3 +25.2 +17.7 .-14.6 .-24.7 0.-5.3 .-21.8 0+7.6 0.-3.5
Table 18: The difference ($\Delta$) in COMET scores on the test sets from English to other languages between our English-German LLM-based DocMT models and their backbones. $\mu_{\Delta}$ indicates the average difference across all the languages in this table.
We also explore the transferability of translation capabilities acquired from one language pair to others. We assess our English-German LLM-based DocMT models on English-to-other-language test sets, comparing their COMET scores to their base models in Table 18. Our results indicate that models with fine-tuned instructions (Llama2-7B and Bloom-7B) consistently exhibit positive transfer effects across all language pairs, while those with instruction-tuned models (Vicuna-7B) benefits only few languages.